# arXiv:2603.02442v1[math.FA] 02 Mar 2026

### 02 Mar 2026 v1

@@ Abstract.
We study Li–Yorke and mean Li–Yorke chaos for weighted composition operators $C_{w,\varphi}$ on Banach spaces of analytic functions on the unit disk $\mathbb{D}$. Under natural conditions on the space, we show that $C_{w,\varphi}$ is (densely) Li–Yorke chaotic if and only if it is not power-bounded, and (densely) mean Li–Yorke chaotic if and only if it is not absolutely Cesàro bounded. These results are applied to Hardy spaces $H^{p}(\mathbb{D})$, $1\leq p\leq\infty$, and weighted Bergman spaces $A^{2}_{\beta}(\mathbb{D})$, $1<\beta<\infty$.
@@ 1. Introduction
The notion of chaos motivating this work was introduced by Li and Yorke [14], who showed that a period-three orbit forces the coexistence of orbits of every period together with an uncountable set of trajectories that neither converge together nor diverge to infinity. This phenomenon, known as Li–Yorke chaos, has since been extended to linear operators on infinite-dimensional Banach and Fréchet spaces [2, 7].
Let $X$ be a Banach space and let $\mathscr{L}(X)$ be the set of all bounded linear operators on $X$. An operator $T\in\mathscr{L}(X)$ is Li–Yorke chaotic if it admits an uncountable scrambled set $S$, meaning that for all $x,y\in S$ with $x\neq y$:
If the set $S$ can be chosen to be dense in $X$, then $T$ is called densely Li–Yorke chaotic. Beyond this notion, several related forms of chaotic behavior have been studied, including distributional chaos, generic chaos, and mean Li–Yorke chaos [2, 6, 7, 3, 4, 5].
A central class of operators is that of weighted composition operators. Given an analytic self-map $\varphi$ defined on the unit disk $\mathbb{D}$, and an analytic weight function $w\colon\mathbb{D}\to\mathbb{C}$, the weighted composition operator $C_{w,\varphi}$ acts on a space of analytic functions by
@@
When $w\equiv 1$, one recovers the classical composition operator $C_{\varphi}$, which has been extensively studied on spaces of analytic functions; see [9, 19]. The present article is devoted to a systematic study of Li–Yorke chaos and mean Li–Yorke chaos for weighted composition operators acting on two families of classical function spaces over the unit disk $\mathbb{D}$: the Hardy spaces $H^{p}(\mathbb{D})$ for $1\leq p\leq\infty$, and the weighted Bergman spaces $A^{2}_{\beta}(\mathbb{D})$ for $1<\beta<\infty$. These spaces provide the natural setting for operator theory on the disk; we refer to [10, 16] for $H^{p}$ and to [11, 13] for weighted Bergman spaces.
@@
The present work addresses this complementary and structurally distinct setting. The key observation is that in spaces of analytic functions such as $H^{p}(\mathbb{D})$ and $A^{2}_{\beta}(\mathbb{D})$, the subspace $H^{\infty}$ plays a dual role that has no analogue in the settings of [4, 5]: it is simultaneously dense in the ambient space and provides sharp norm estimates for the iterates of $C_{w,\varphi}$ in terms of the weight iterates $w^{(n)}$. This structure is captured by two conditions (C.1) and (C.2) introduced in Section 3, which serve as the foundation of our approach. Under these conditions, we prove that $C_{w,\varphi}$ is (densely) Li–Yorke chaotic if and only if it is not power-bounded, and (densely) mean Li–Yorke chaotic if and only if it is not absolutely Cesàro bounded. These abstract results are then applied to obtain complete characterizations of Li–Yorke chaos and mean Li–Yorke chaos for weighted composition operators on $H^{p}(\mathbb{D})$, $1\leq p<\infty$, and on $A^{2}_{\beta}(\mathbb{D})$, $1<\beta<\infty$, expressed entirely in terms of the symbols $w$ and $\varphi$. The case $p=\infty$ is treated separately in Section 4, where the lack of separability of $H^{\infty}$ requires an adapted argument. To illustrate the sharpness of our conditions, Section 5 presents a concrete example of a weighted composition operator that is Li–Yorke chaotic on $H^{p}$ for $1\leq p\leq 2$.
@@ 2. Li–Yorke Chaotic, Power Bounded, and Absolutely Cesàro Bounded Operators
In this section, we introduce the essential results about the different kinds of Li–Yorke chaos, as well as the notions of power-bounded and absolutely Cesàro bounded operators. Before turning to the main results of this section, we briefly recall the definition of residual set and we introduce an important result for operators defined on separable Banach spaces.
@@ 2.1. Notions of Li–Yorke chaos
Previously, we provided a definition of Li–Yorke chaos. Recall that a vector $x\in X$ is said to be irregular for $T$ if
@@
By [7, Theorem 9], an operator is Li–Yorke chaotic if and only if it admits an irregular vector.
@@ 2.3. Li–Yorke chaotic operators are not power-bounded
Conversely, suppose that $T$ is not Li-Yorke chaotic. Let $x\in X$ whose orbit under $T$ has a subsequence converging to zero; this implies that $\lim_{n\to\infty}\inf\|T^{n}x\|=0$. Since $T$ has no irregular vectors, it follows that $\limsup_{n\to\infty}\|T^{n}x\|=0$ as well, and therefore $\lim_{n\to\infty}T^{n}x=0$. In this setting, the set $\mathrm{NS}(T)$ defines a subspace. If $\mathscr{X}$ is finite-dimensional, $\mathscr{X}=\mathrm{NS}(T)$, thus $T$ is power-bounded in $\mathscr{X}$. Otherwise, $\mathrm{NS}(T)$ is dense and by Proposition 1, we conclude that $\mathrm{NS}(T)$ is residual in $\mathscr{X}$. Now define $\mathscr{R}$ as the set of all vectors $x\in\mathscr{X}$ whose orbit under $T$ is unbounded. If $T$ was not power-bounded in $\mathscr{X}$, by the Banach–Steinhaus [17, Theorem 2.5], $\mathscr{X}\setminus\mathscr{R}$ is a proper subset of first category on $X$, hence $\mathscr{R}$ is residual in $\mathscr{X}$. Since the intersection $\mathrm{NS}(T)\cap\mathscr{R}$ is residual in $\mathscr{X}$ (and thus dense in $\mathscr{X}$), any element $y\in\mathrm{NS}(T)\cap\mathscr{R}$ would be an irregular vector. Since we are assuming that $T$ is not Li-Yorke chaotic, we conclude that is power bounded $\mathscr{X}$. ∎
@@
If (iii) holds, proceeding as the proof of Proposition 2, one can show that the set $\mathscr{R}_{2}$ of all vectors $x\in X$ whose orbit under $T$ is unbounded is residual. Moreover, by the hypothesis, the set $\mathscr{R}_{1}$ of all vectors $x\in X$ whose orbit under $T$ has a subsequence converging to zero is dense in $X$, and therefore residual by Proposition 1; then $\mathscr{R}_{1}\cap\mathscr{R}_{2}$ is residual and every vector is irregular, which implies (ii). $\text{(ii)}\Rightarrow\text{(i)}$ follows by [7, Theorem 10], while $\text{(i)}\Rightarrow\text{(iii)}$ follows once again from Proposition 2. ∎
3. Li–Yorke chaos for weighted composition operators
@@ 3.1. Setup and structural conditions
C.1 $H^{\infty}$ is dense in $X$. C.2 Every weighted composition operator on $X$, $C_{w,\varphi}$, where $\varphi:\mathbb{D}\to\mathbb{D}$ is an analytic function and $w\in H^{\infty}$, is bounded and satisfies $\|(C_{w,\varphi})^{n}f\|_{X}\leq\|w^{(n)}\|_{X}\|f\|_{\infty},\quad\forall f\in H^{\infty},\quad\forall n\in\mathbb{N}.$
@@
  • i)
    $\lim_{k\to\infty}\dfrac{1}{N_{k}}\sum_{j=1}^{N_{k}}\|w^{(j)}\|_{X}=0$;
  • ii)
    $\sup_{N\in\mathbb{N}}\dfrac{1}{N}\sum_{j=1}^{N}\|w^{(j)}\|_{X}=\infty.$
@@ 3.2. The Hardy spaces $H^{p}$ for $1\leq p<\infty$
where $m$ is normalized Lebesgue measure on $\mathbb{T}$. The $H^{p}$ spaces are separable Banach spaces for $p\geq 1$. From now on $p$ will denote a real number $p\in[1,\infty)$.
Let $C_{w,\varphi}$ be a weighted composition operator on $H^{p}$. If the weight function $w$ belongs to $H^{\infty}$, then $C_{w,\varphi}$ is a bounded operator on $H^{p}$; see [15, 18] for references.
@@
For any weighted composition operator $C_{w,\varphi}$ on $H^{p}$:
@@
3.3. The weighted Bergman spaces $A^{2}_{\beta}$ for $1<\beta<\infty$
Let $dA(z)$ be the normalized area measure on $\mathbb{D}$, and $1<\beta<\infty$. The weighted Bergman space $A^{2}_{\beta}(\mathbb{D})$ is the space of all analytic functions in $\mathbb{D}$ such that the norm
$\|f\|^{2}_{\beta}:=\int_{\mathbb{D}}|f(z)|^{2}\,dA_{\beta}(z)<\infty,$
where $dA_{\beta}(z)=(\beta+1)(1-|z|^{2})^{\beta}\,dA(z)$. The weighted Bergman space $A^{2}_{\beta}$ is a separable Hilbert space with the inner product
@@ 3.3. The weighted Bergman spaces $A^{2}_{\beta}$ for $1<\beta<\infty$
where $(\widehat{f}(n))_{n\in\mathbb{N}}$ and $(\widehat{g}(n))_{n\in\mathbb{N}}$ are the sequences of Maclaurin coefficients for $f$ and $g$ respectively, and $\Gamma$ is the Gamma function. It is straightforward to verify that $H^{\infty}$ is a subspace of $A^{2}_{\beta}$ and $\|f\|_{\beta}\leq\|f\|_{\infty}$.
Let $C_{w,\varphi}$ be the weighted composition operator on $A^{2}_{\beta}$. If $w\in H^{\infty}$, $C_{w,\varphi}$ defines a bounded operator since the operator $M_{\psi}:A^{2}_{\beta}\rightarrow A^{2}_{\beta}$ by $(M_{\psi}f)(z)=\psi(z)f(z)$ is bounded and the composition operator $C_{\varphi}$ as well, see [11, Page 532].
The next Lemma shows that condition (C.2) holds for $X=A^{2}_{\beta}$.
@@
For any weighted composition operator $C_{w,\varphi}$ on $A^{2}_{\beta}$:
$\|(C_{w,\varphi})^{n}f\|_{\beta}\leq\|w^{(n)}\|_{\beta}\|f\|_{\infty},\quad\forall f\in H^{\infty},\quad\forall n\in\mathbb{N}.$
@@
$\|w^{(n)}\cdot(f\circ\varphi^{n})\|^{2}_{\beta}$ $=\int_{\mathbb{D}}|w^{(n)}(z)(f\circ\varphi^{n})(z)|^{2}\,dA_{\beta}(z)$ $\leq\|f\|_{\infty}^{2}\int_{\mathbb{D}}|w^{(n)}(z)|^{2}\ dA_{\beta}(z)$
$\leq\|w^{(n)}\|^{2}_{\beta}\|f\|_{\infty}^{2}.$
Therefore, $\|(C_{w,\varphi})^{n}f\|_{\beta}\leq\|w^{(n)}\|_{\beta}\|f\|_{\infty},\ \ \forall f\in H^{\infty}.$
It is worth noting that the closure of $H^{\infty}$ in the $A^{2}_{\beta}$-norm topology is the whole space $A^{2}_{\beta}$. Indeed, since $\operatorname{span}\{z^{n}\}_{n\geq 0}\subset H^{\infty}\quad\text{and}\quad A^{2}_{\beta}=\overline{\operatorname{span}\{z^{n}\}_{n\geq 0}}^{A^{2}_{\beta}},$ (see [11, Page 532]) it follows that $\overline{H^{\infty}}^{A^{2}_{\beta}}=A^{2}_{\beta}$. Therefore, condition (C.1) holds.
Note that if $\varphi:\mathbb{D}\to\mathbb{D}$ is an analytic self-map and $w\in H^{\infty}$, then the weighted composition operator $C_{w,\varphi}$ is bounded on both $H^{2}$ and $A^{2}_{\beta}$.
@@
    $\sup\{\|w^{(n)}\|_{\beta}:n\in\mathbb{N}\}=\infty$
@@
According to [11], $\|f\|_{\beta}\leq C\|f\|_{2},$ for all $f\in H^{2}.$ Then the result follows by Corollary 1 for $X=A^{2}_{\beta}$ and $X=H^{2}$, respectively. ∎
@@
  • i)
    $\lim_{k\to\infty}\dfrac{1}{N_{k}}\sum_{j=1}^{N_{k}}\|w^{(j)}\|_{2}=0$, and
  • ii)
    $\sup_{N\in\mathbb{N}}\dfrac{1}{N}\sum_{j=1}^{N}\|w^{(j)}\|_{\beta}=\infty.$
@@
According to [11], $\|f\|_{\beta}\leq C\|f\|_{2},$ for all $f\in H^{2}.$ Then the result follows by Corollary 2 for $X=A^{2}_{\beta}$ and $X=H^{2}$, respectively. ∎

### 13 Mar 2026 v2

# Li–Yorke chaotic weighted composition operators on Hardy and weighted Bergman spaces over the unit disk

@@ Abstract.
We study Li–Yorke and mean Li–Yorke chaos for weighted composition operators $C_{w,\varphi}$ on Banach spaces of analytic functions on the unit disk $\mathbb{D}$. Under natural conditions on the space, we show that $C_{w,\varphi}$ is (densely) Li–Yorke chaotic if and only if it is not power-bounded, and (densely) mean Li–Yorke chaotic if and only if it is not absolutely Cesàro bounded. These results are applied to Hardy spaces $H^{p}(\mathbb{D})$, $1\leq p\leq\infty$, and weighted Bergman spaces $A^{p}_{\beta}(\mathbb{D})$, $1<\beta<\infty$ and $1<p<\infty$.
@@ 1. Introduction
The notion of chaos motivating this work was introduced by Li and Yorke [16], who showed that a period-three orbit forces the coexistence of orbits of every period together with an uncountable set of trajectories that neither converge together nor diverge to infinity. This phenomenon, known as Li–Yorke chaos, captures the coexistence of pairs of orbits that are recurrently close yet persistently apart; it has since been extended to linear operators on infinite-dimensional Banach and Fréchet spaces [3, 8]. Beyond this notion, several related forms of chaotic behavior have been studied, including distributional chaos, generic chaos, and mean Li–Yorke chaos [3, 7, 8, 4, 5, 6].
Weighted composition operators constitute a central and well-studied class in operator theory. Given an analytic self-map $\varphi\colon\mathbb{D}\to\mathbb{D}$ and an analytic weight function $w\colon\mathbb{D}\to\mathbb{C}$, the weighted composition operator $C_{w,\varphi}$ acts on a space of analytic functions by
@@
When $w\equiv 1$, one recovers the classical composition operator $C_{\varphi}$, which has been extensively studied on spaces of analytic functions; see [11, 20]. The present article is devoted to a systematic study of Li–Yorke chaos and mean Li–Yorke chaos for weighted composition operators acting on two families of classical function spaces over the unit disk $\mathbb{D}$: the Hardy spaces $H^{p}(\mathbb{D})$ for $1\leq p\leq\infty$, and the weighted Bergman spaces $A^{p}_{\beta}(\mathbb{D})$ for $1<\beta<\infty$, $1<p<\infty$. These spaces provide the natural setting for operator theory on the disk; we refer to [12, 17] for $H^{p}$ and to [13, 15] for weighted Bergman spaces.
@@
The present work addresses this complementary and structurally distinct setting. The key observation is that in spaces of analytic functions such as $H^{p}(\mathbb{D})$ and $A^{p}_{\beta}(\mathbb{D})$, the subspace $H^{\infty}$ plays a dual role that has no analogue in the settings of [5, 6]: it is simultaneously dense in the ambient space and provides sharp norm estimates for the iterates of $C_{w,\varphi}$ in terms of the weight iterates $w^{(n)}$. This structure is captured by two conditions (C.1) and (C.2) introduced in Section 3, which serve as the foundation of our approach. Under these conditions, we prove that $C_{w,\varphi}$ is (densely) Li–Yorke chaotic if and only if it is not power-bounded, and (densely) mean Li–Yorke chaotic if and only if it is not absolutely Cesàro bounded. These abstract results are then applied to obtain sufficient conditions for Li–Yorke chaos and mean Li–Yorke chaos for weighted composition operators on $H^{p}(\mathbb{D})$, $1\leq p<\infty$, and on $A^{p}_{\beta}(\mathbb{D})$, $1<\beta<\infty$, $1<p<\infty$, expressed entirely in terms of the symbols $w$ and $\varphi$. The case $p=\infty$ is treated separately in Section 4, where the lack of separability of $H^{\infty}$ requires an adapted argument. To illustrate the sharpness of our conditions, Section 5 presents examples of weighted composition operators that are Li–Yorke chaotic on $A^{2}_{\beta}(\mathbb{D})$ and $H^{p}(\mathbb{D})$ for $1\leq p\leq 2$. One of these examples additionally shows that the hypotheses of Theorem 4 are not necessary for Li–Yorke chaos to occur.
@@ 2. Li–Yorke Chaotic, Power Bounded, and Absolutely Cesàro Bounded Operators
Let $X$ be a Banach space and let $\mathscr{L}(X)$ be the set of all bounded linear operators on $X$. In this section, we introduce the essential results about the different kinds of Li–Yorke chaos, as well as the notions of power-bounded and absolutely Cesàro bounded operators. Before turning to the main results of this section, we briefly recall the definition of residual set and we introduce an important result for operators defined on separable Banach spaces.
@@ 2.1. Notions of Li–Yorke chaos
An operator $T\in\mathscr{L}(X)$ is Li–Yorke chaotic if it admits an uncountable scrambled set $S$, meaning that for all $x,y\in S$ with $x\neq y$:
If the set $S$ can be chosen to be dense in $X$, then $T$ is called densely Li–Yorke chaotic.
Recall that a vector $x\in X$ is said to be irregular for $T$ if
@@
By [8, Theorem 9], an operator is Li–Yorke chaotic if and only if it admits an irregular vector.
@@ 2.3. Li–Yorke chaotic operators are not power-bounded
Conversely, suppose that $T$ is not Li-Yorke chaotic. Let $x\in X$ whose orbit under $T$ has a subsequence converging to zero; this implies that $\lim_{n\to\infty}\inf\|T^{n}x\|=0$. Since $T$ has no irregular vectors, it follows that $\limsup_{n\to\infty}\|T^{n}x\|=0$ as well, and therefore $\lim_{n\to\infty}T^{n}x=0$. In this setting, the set $\mathrm{NS}(T)$ defines a subspace. If $\mathscr{X}$ is finite-dimensional, $\mathscr{X}=\mathrm{NS}(T)$, thus $T$ is power-bounded in $\mathscr{X}$. Otherwise, $\mathrm{NS}(T)$ is dense and by Proposition 1, we conclude that $\mathrm{NS}(T)$ is residual in $\mathscr{X}$. Now define $\mathscr{R}$ as the set of all vectors $x\in\mathscr{X}$ whose orbit under $T$ is unbounded. If $T$ was not power-bounded in $\mathscr{X}$, by the Banach–Steinhaus [18, Theorem 2.5], $\mathscr{X}\setminus\mathscr{R}$ is a proper subset of first category on $X$, hence $\mathscr{R}$ is residual in $\mathscr{X}$. Since the intersection $\mathrm{NS}(T)\cap\mathscr{R}$ is residual in $\mathscr{X}$ (and thus dense in $\mathscr{X}$), any element $y\in\mathrm{NS}(T)\cap\mathscr{R}$ would be an irregular vector. Since we are assuming that $T$ is not Li-Yorke chaotic, we conclude that is power bounded $\mathscr{X}$. ∎
@@
If (iii) holds, proceeding as the proof of Proposition 2, one can show that the set $\mathscr{R}_{2}$ of all vectors $x\in X$ whose orbit under $T$ is unbounded is residual. Moreover, by the hypothesis, the set $\mathscr{R}_{1}$ of all vectors $x\in X$ whose orbit under $T$ has a subsequence converging to zero is dense in $X$, and therefore residual by Proposition 1; then $\mathscr{R}_{1}\cap\mathscr{R}_{2}$ is residual and every vector is irregular, which implies (ii). $\text{(ii)}\Rightarrow\text{(i)}$ follows by [8, Theorem 10], while $\text{(i)}\Rightarrow\text{(iii)}$ follows once again from Proposition 2. ∎
3. Li–Yorke chaos for weighted composition operators on spaces of analytic functions
@@ 3.1. Setup and structural conditions
C.1 $H^{\infty}$ is dense in $X$. C.2 Every weighted composition operator $C_{w,\varphi}$ on $X$, where $\varphi:\mathbb{D}\to\mathbb{D}$ is an analytic self-map and $w\in H^{\infty}$, is bounded and satisfies $\|(C_{w,\varphi})^{n}f\|_{X}\leq\|w^{(n)}\|_{X}\|f\|_{\infty},\quad\forall f\in H^{\infty},\quad\forall n\in\mathbb{N}.$
@@
$\lim_{k\to\infty}\dfrac{1}{N_{k}}\sum_{j=1}^{N_{k}}\|w^{(j)}\|_{X}=0\quad\text{and}\quad\displaystyle\sup_{N\in\mathbb{N}}\dfrac{1}{N}\sum_{j=1}^{N}\|w^{(j)}\|_{X}=\infty.$
@@ 3.2. The Hardy spaces $H^{p}$ for $1\leq p<\infty$
where $m$ is normalized Lebesgue measure on $\mathbb{T}$. The $H^{p}$ spaces are separable Banach spaces for $1\leq p<\infty$.
The space $H^{2}$ is a Hilbert space with inner product
$\langle f,g\rangle=\sum_{n=0}^{\infty}a_{n}\overline{b_{n}},$
where $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ are the Maclaurin coefficients for $f$ and $g$ respectively.
If the weight function $w$ belongs to $H^{\infty}$ and $\varphi$ is an analytic self-map on $\mathbb{D}$, it is known that $\varphi$ induces a bounded composition operator on $H^{p}$ (see [19]) and the multiplication operator induced by $w$ is also bounded on $H^{p}$, then $C_{w,\varphi}$ is a bounded operator on $H^{p}$.
@@
For $1\leq p<\infty$, let $C_{w,\varphi}$ be a weighted composition operator on $H^{p}$. Then
@@
3.3. The weighted Bergman spaces $A^{p}_{\beta}$ for $1<\beta<\infty$ and $1<p<\infty$
Let $dA(z)$ be the normalized area measure on $\mathbb{D}$, and $1<\beta<\infty$. The weighted Bergman space $A^{p}_{\beta}(\mathbb{D})$ is the space of all analytic functions in $\mathbb{D}$ such that the norm
$\|f\|^{p}_{\beta_{p}}:=\int_{\mathbb{D}}|f(z)|^{p}\,dA_{\beta}(z)<\infty,$
where $dA_{\beta}(z)=(\beta+1)(1-|z|^{2})^{\beta}\,dA(z)$. The weighted Bergman space $A^{p}_{\beta}$ is a separable infinite-dimensional Banach space. It is straightforward to verify that $H^{\infty}$ is a subspace of $A^{p}_{\beta}$ and $\|f\|_{\beta_{p}}\leq C\|f\|_{\infty}$ for some $C>0$.
The weighted Bergman space $A^{2}_{\beta}$ is a Hilbert space with the inner product
@@ 3.3. The weighted Bergman spaces $A^{p}_{\beta}$ for $1<\beta<\infty$ and $1<p<\infty$
where $(\widehat{f}(n))_{n\in\mathbb{N}}$ and $(\widehat{g}(n))_{n\in\mathbb{N}}$ are the sequences of Maclaurin coefficients for $f$ and $g$ respectively, and $\Gamma$ is the Gamma function.
If the weight function $w$ belongs to $H^{\infty}$ and $\varphi$ is an analytic self-map on $\mathbb{D}$, it is known that $\varphi$ induces a bounded composition operator on $A^{p}_{\beta}$ (see [9, Proposition 7.1]) and the multiplication operator induced by $w$ is also bounded on $A^{p}_{\beta}$, then $C_{w,\varphi}$ is a bounded operator on $A^{p}_{\beta}$.
The next Lemma shows that condition (C.2) holds for $X=A^{p}_{\beta}$.
@@
For $1<p<\infty$, let $C_{w,\varphi}$ be a weighted composition operator on $A^{p}_{\beta}$. Then
$\|(C_{w,\varphi})^{n}f\|_{\beta_{p}}\leq\|w^{(n)}\|_{\beta_{p}}\|f\|_{\infty},\quad\forall f\in H^{\infty},\quad\forall n\in\mathbb{N}.$
@@
$\|w^{(n)}\cdot(f\circ\varphi^{n})\|^{p}_{\beta_{p}}$ $=\int_{\mathbb{D}}|w^{(n)}(z)(f\circ\varphi^{n})(z)|^{p}\,dA_{\beta}(z)$ $\leq\|f\|_{\infty}^{p}\int_{\mathbb{D}}|w^{(n)}(z)|^{p}\ dA_{\beta}(z)$
$\leq\|w^{(n)}\|^{p}_{\beta_{p}}\|f\|_{\infty}^{p}.$
Therefore, $\|(C_{w,\varphi})^{n}f\|_{\beta_{p}}\leq\|w^{(n)}\|_{\beta_{p}}\|f\|_{\infty},\ \ \forall f\in H^{\infty}.$
It is worth noting that the closure of $H^{\infty}$ in the $A^{p}_{\beta}$-norm topology is the whole space $A^{p}_{\beta}$. Indeed, since $\operatorname{span}\{z^{n}\}_{n\geq 0}\subset H^{\infty}$ and the polinomials are dense in $A^{p}_{\beta}$ it follows that $\overline{H^{\infty}}^{A^{p}_{\beta}}=A^{p}_{\beta}$ (see [15] for a reference). Therefore, condition (C.1) holds.
@@
    $\sup\{\|w^{(n)}\|_{\beta_{2}}:n\in\mathbb{N}\}=\infty$
@@
According to [13], we have $\|f\|_{\beta}\leq C\|f\|_{2}$, for all $f\in H^{2}$. Then the result follows by Corollary 1 for $X=A^{2}_{\beta}$ and $X=H^{2}$, respectively. ∎
@@
$\lim_{k\to\infty}\dfrac{1}{N_{k}}\sum_{j=1}^{N_{k}}\|w^{(j)}\|_{2}=0\quad\text{and}\quad\displaystyle\sup_{N\in\mathbb{N}}\dfrac{1}{N}\sum_{j=1}^{N}\|w^{(j)}\|_{\beta_{2}}=\infty.$
@@
According to [13], we have $\|f\|_{\beta}\leq C\|f\|_{2}$, for all $f\in H^{2}$. Then the result follows by Corollary 2 for $X=A^{2}_{\beta}$ and $X=H^{2}$, respectively. ∎